Grade 11Algebra

f(x)= 9x/(9x +3)
then find the value of
f(1/2008)+f(2/2008)+f(3/2008)+.......+f(2007/2008)=?

Nikunj
9 Years agoGrade 11

1 Answer

Rabi Islam
9 Years ago
f (x) = 9^x/(9^x+3) and f(1-x) =3/(9^x+3)f (2007/2008) = f (1-1/2008)Similarly next allf (x)+f (1-x)=1f (1/2008)+f (2/2008)+.....f (1-2/2008)+f (1-1/2008)making pairstoral pairs, 2007/2 =1003.5 (if f (x)=0.5 and f (1-x)=0.5 , Then f (x)+f (1-x)=1)So total pairs , 1+1+1+....+1003+0.5 = 1003.5